OpenStudy (anonymous):

The measures of two sides of a triangle are given. If P is the perimeter and P=10x+5y find the measure of the third side. Side one: 3x+4y Side two: 5x-y Side three: looking for

OpenStudy (anonymous):

... 10x+5y-3x-4y-5x-y=2x-4y

OpenStudy (anonymous):

(3x+4y)+(5x-y)=8x+3y (10x+5y)-(8x+3y)=2x +2y Side three is: 2x +2y